OpenStudy (anonymous):

the sum of a number and twice its square is 6. Find the number

hartnn (hartnn):

let the number be 'x' sum of a number and twice its square = x+2x^2 so, 2x^2 + x=6 so, 2x^2 +x-6=0 can you solve this quadratic by using the formula i gave you ?

OpenStudy (anonymous):

Okay wait

hartnn (hartnn):

do you want me to post that formula again ?

OpenStudy (anonymous):

2 and -3/2

OpenStudy (anonymous):

Nope i got the formula :)

OpenStudy (anonymous):

-3/2 and 2 is this right?

OpenStudy (phi):

the signs are the other way around: -2 and 3/2

hartnn (hartnn):

isn't it -2 and +3/2 ??

OpenStudy (anonymous):

Oh i'm sorry i forgot the negative sign in 1

hartnn (hartnn):

if that number is integer , then only -2 is the solution, but if its real number then both -2 and 3/2 are answers.

OpenStudy (anonymous):

Thanks again!

hartnn (hartnn):

welcome again!